Fermi National Laboratory
 

Accelerator Update for September 27, 2006 - September 29, 2006

 

Wednesday September 27

The day shift began with Operations monitoring a Recycler stash, with the Tevatron (TeV) recovering from a quench, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for colliding beam, with the Antiproton Source (Pbar) stacking, and with beam to the Linac, Booster, Main Injector (MI), Recycler, NuMI and MiniBooNE.

Operations reported at 8:24 AM that the quench was due to a reversing switch fault on C:DQ0. Experts began repairs at 8:40 AM.

At 10:24 AM, Operations received a low cooling level alarm on a TeV kicker (F17K2).  It tripped off four minutes later.

The Recycler completed a stash at 11:21 AM.

The Recycler completed a stash at 3:05 PM.

At 3:26 PM, Operations reported that a Debuncher RF station (DRF2) wasn’t putting out any power. Operations contacted an expert.

At 5:48 PM, Operations began preparing the TeV F sector for an access to investigate F17K2.

Operations put the TeV through a dry squeeze at 7:16 PM.

Operations put the TeV into shot setup at 7:51 PM.

At 8:38 PM, the TeV suffered a quench due to a low beta magnet power supply (AQ7).  Operations called in an expert.

Operations put the TeV through a dry squeeze at 9:49 PM.

Operations put the TeV into shot setup at 10:25 PM.

Thursday September 28

The midnight shift began with Operations monitoring the stash and stack, with the TeV in shot setup, and with NuMI and MiniBooNE taking beam.

Operations established store 4981 around 12:05 AM with an initial luminosity of 90.59E30.

The Recycler completed a stash at 1:06 AM.

At 4:15 AM, the TeV quenched at sector F17.  Operators discovered that a late stacking pulse was the problem.

Operations prepared D0 and CDF at 5:37 AM for access to their collision halls.

By 6:42 AM, Operations had begun preparing Booster, NuMI, MiniBooNE, and the TeV for access.

The Neutron Therapy Facility (NTF) reported at 12:37 PM that they had finished patient treatment for the day.

Experts prepared Linac for safety system testing at 12:40 PM. The testing was completed by 2:17 PM.

Operations established beam to the Linac dump at 3:37 PM.

Technicians completed the TeV sector F17 kicker magnet replacement and were out of the enclosure by 11:41 PM.

Operations established beam to MiniBooNE at 11:52 PM.

Friday September 29

The midnight shift began with Operations monitoring the stash and stack, and with MiniBooNE taking beam.

At 1:20 AM, Pbar experts working on the lithium lens reported a power supply problem.  Operations called in an expert.

At 1:31 AM, Operations reported that they could not close the MI-20 VCB.  Experts came in to investigate.

Operations put the TeV through a dry squeeze at 2:19 AM.

Operators were able to close the MI-20 VCB at 3:58 AM.  Experts repaired a bad relay.

Operations established beam to NuMI at 4:16 AM.

Operations put the TeV through a wet squeeze at 5:05 AM.

Operations put the TeV into shot setup at 5:37 AM.

Pbar experts reported at 7:06 AM that they had completed all repairs on the lithium lens and power supply.

At 7:09 AM, the TeV suffered a quench at sector E48.

The Plans for Friday and the Weekend

The plans for today are to find the cause of the TeV quench. Hopefully, a new store will get established by mid afternoon. The Run Coordinator will allow studies that have no impact on stacking and stashing during the quench investigations. The weekend will consist of stacking, stashing, and storing.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 09/29/2006    email Fermilab